# Is My Data in Your Retrieval Database? Membership Inference Attacks Against Retrieval Augmented Generation

### v1

@@ Abstract
This study addresses this gap by introducing an efficient and easy-to-use method for conducting MIA against RAG systems. We demonstrate the effectiveness of our attack using two benchmark datasets and multiple generative models, showing that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt in both black-box and gray-box settings. Our findings highlight the importance of implementing security countermeasures in deployed RAG systems to protect the privacy and security of retrieval databases.
@@ 1 Introduction
In this study, we introduce a method that is both efficient and easy to use for conducting MIAs against RAG systems, with the objective of determining whether a specific data sample is present in the retrieval database. We assess the effectiveness of our attack using two benchmark datasets that represent scenarios where privacy may be a concern, and employing multiple generative models.
The findings of our study reveal that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt, underscoring the importance of implementing security countermeasures in deployed RAG systems.
@@ 2.2 Membership Inference Attacks
Membership inference attacks [15, 9] are a type of privacy threat where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model. This carries significant privacy implications as it can potentially reveal sensitive information about individuals, even if the model does not directly release any personal data.
Formally, an attacker aims to determine the membership of a sample $x$ in the training data of a target model $\mathcal{D}_{m}$, i.e., to check if $x\in\mathcal{D}_{m}$, which is known as sample-level membership inference. Typically, these attacks involve calculating one or more metrics on the target model’s outputs that reflect the probability of the sample being a part of the training set, such as the model outputs’ entropy or log-probabilities [3]. Several metrics may be computed for each sample and then fused together using a machine learning model, known as an attack model, which in turn outputs the probability of a sample being a member of the training set. Additionally, an attacker may also aim to determine the membership of a certain user, i.e., to check if a user’s data is part of the training set, which is known as user-level membership inference [14]. However throughout this paper we will address the membership inference challenge from a sample-level perspective.
@@ 2.3 Threat Model
This paper considers a black-box scenario, in which the attacker has access solely to the user prompt and the resulting generated output from the RAG system. The attacker can modify the user prompt in any manner they deem appropriate; however, they possess no knowledge of the underlying models $E$ or $G$, nor the prompt templates that are being used by these models. Furthermore, the attacker has no information regarding the deployment details, such as the type of retrieval database employed.
@@ 3 Methodology
In this section we present our membership inference attack.
The input to the generative model used by the system consists of the RAG tempate into which the retrieved database entries and the user prompt are inserted, as depicted in Figure 1. Since the attacker’s interaction with the system is limited to the user prompt part of the RAG template, the attack must be based on a prompt that achieves both of the following goals:
@@ 3.1 Experimental Setup
Throughout the evaluation we used two datasets that represent practical scenarios, in which privacy can be critical:
@@
[…]
@@ 4 Results
4.1 Impact of Attack Prompt
@@
Unsurprisingly, the results in the black-box setting are inferior to those in the gray-box setting. This is especially the case for the flan model, with an improvement of up to 25% in the gray-box setting. This means that the log-likelihood values for member documents are significantly higher than for non-member documents, indicating a higher confidence of the model in its response.
However, when looking at the llama and mistral models, the average difference is only up to 12%, and in some cases even lower, depending on the prompt.
To further explore this difference, we analyze the percentage of samples that are correctly retrieved from the database for each prompt. We found that over 95% of the member samples are indeed retrieved for both datasets. This is in contrast with the non-member samples, which are retrieved in nearly 0% of the cases. The full results of this analysis can be found in Appendix B.
Thus, we conclude that the flan model is more grounded to the content of its input prompt, and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
4.2 Results Summary
To quantify the overall MIA risk, we summarize the best results (across prompts) for each combination of model and dataset in Table 2 . We can see that, on average, the attack success rate is highest for flan, in both threat models.
In addition, we observe that the overall risk in the black-box setting is similar between almost all models, with the exception of the HealthCareMagic-llama case. However in the gray-box setting the results vary more.
Our lowest attack AUC-ROC score of 0.74 is quite high compared to previous research on sample-level MIA in language models [5, 14]. Overall, our black-box attack achieves an average AUC-ROC of 0.80 and the gray-box attack achieves an average AUC-ROC of 0.90. The best gray-box attack on flan accomplishes perfect performance. These results highlight the risk associated with deploying RAG based systems without proper defense mechanisms.
@@ 5 Conclusion
Our attack achieves a very high average AUC-ROC of 0.90 and 0.80 in the gray-box and the black-box threat models, respectively, and for some models almost perfect performance. This underscores the need to develop defenses and countermeasures for this attack scenario. We hope that the research community will proceed to explore the risk of membership inference in RAG-based systems and employ the ideas from this paper as a baseline.
@@ B Detailed Results of Retrieval Matches
Tables 4 and 5 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively, for the different attack prompts. We see that the chosen attack prompts do not differ in their influence on the retrieval accuracy.
@@
C Model outputs without a clear yes/no
answer
The number of model outputs that did not contain either of the ”Yes” or ”No” tokens are shown in Table 6.

### v2

@@ Abstract
This study addresses this gap by introducing an efficient and easy-to-use method for conducting MIA against RAG systems. We demonstrate the effectiveness of our attack using two benchmark datasets and multiple generative models, showing that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt in both black-box and gray-box settings. Moreover, we introduce an initial defense strategy based on adding instructions to the RAG template, which shows high effectiveness for some datasets and models. Our findings highlight the importance of implementing security countermeasures in deployed RAG systems and developing more advanced defenses to protect the privacy and security of retrieval databases.
Keywords:
AI Privacy; Membership Inference; RAG; Large Language Models
@@ 1 Introduction
In this study, we introduce a method that is both efficient and easy to use for conducting MIAs against RAG systems, with the objective of determining whether a specific data sample is present in the retrieval database. We assess the effectiveness of our attack using two benchmark datasets that represent scenarios where privacy may be a concern, and employing multiple generative models. Moreover, we introduce an initial defense strategy based on adding instructions to the RAG template to prevent the model from responding to the attack prompt. This approach seems to be highly effective for some datasets and models, but should be further developed to provide a more comprehensive solution.
The findings of our study reveal that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt, underscoring the importance of developing appropriate defenses and implementing security countermeasures in deployed RAG systems.
@@ 2.2 Membership Inference Attacks
Membership inference attacks [15, 9] are a type of privacy threat, where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model. This carries significant privacy implications as it can potentially reveal sensitive information about individuals, even if the model does not directly release any personal data.
Formally, an attacker aims to determine the membership of a sample $x$ in the training data of a target model $\mathcal{D}_{m}$, i.e., to check if $x\in\mathcal{D}_{m}$, which is known as sample-level membership inference. Typically, these attacks involve calculating one or more metrics on the target model’s outputs that reflect the probability of the sample being a part of the training set, such as the model outputs’ entropy or log-probabilities [3]. Several metrics may be computed for each sample and then fused together using a machine learning model, known as an attack model, which in turn outputs the probability of a sample being a member of the training set. Additionally, an attacker may also aim to determine the membership of a certain user, i.e., to check if a user’s data is part of the training set, which is known as user-level membership inference [14]. Throughout this paper we will address the membership inference challenge from a sample-level perspective.
@@ 2.3 Threat Model
This paper considers a black-box scenario in which the attacker has access solely to the user prompt and the resulting generated output from the RAG system. The attacker can modify the user prompt in any manner they deem appropriate; however, they possess no knowledge of the underlying models $E$ or $G$, nor the prompt templates that are being used by these models. Furthermore, the attacker has no information regarding the deployment details, such as the type of retrieval database employed.
@@ 3 Methodology
In this section we present our membership inference attack. The input to the generative model used by the system consists of the RAG tempate into which the retrieved database entries and the user prompt are inserted, as depicted in Figure 1. Since the attacker’s interaction with the system is limited to the user prompt part of the RAG template, the attack must be based on a prompt that achieves both of the following goals:
@@ 3.1 Experimental Setup
Throughout the evaluation we used two datasets that represent practical scenarios in which privacy can be critical:
@@
Please answer the question using the context provided.
If the question is unanswerable, say unanswerable”.
Question: {user prompt}.
Context:
{context}
Question: {user prompt}
@@ 4 Results
4.1 Impact of Attack Prompts
@@
[…]
Unsurprisingly, the results in the black-box setting are inferior to those in the gray-box setting. This is especially the case for the flan model, with an improvement of up to 25% in the gray-box setting. This means that the log-likelihood values for member documents are significantly higher than for non-member documents, indicating a higher confidence of the model in its response. However, when looking at the llama and mistral models, the average difference is only up to 12%, and in some cases even lower, depending on the prompt.
To further explore this difference, we analyze the percentage of samples that are correctly retrieved from the database for each prompt. We found that over 95% of the member samples are indeed retrieved for both datasets. This is in contrast with the non-member samples, which are retrieved in nearly 0% of the cases. The full results of this analysis can be found in Appendix B. Thus, we conclude that the flan model is more grounded to the content of its input prompt (context grounding), and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
4.2 Attack Results Summary
To quantify the overall MIA risk, we summarize the best results (across prompts) for each combination of model and dataset in Table 2 . We can see that, on average, the attack success rate is highest for flan, in both threat models. In addition, we observe that the overall risk in the black-box setting is similar between almost all models, with the exception of the HealthCareMagic-llama case. However in the gray-box setting the results vary more.
@@ 4.2 Attack Results Summary
Our lowest attack AUC-ROC score of 0.74 is quite high compared to previous research on sample-level MIA in language models [5, 14]. Overall, our black-box attack achieves an average AUC-ROC of 0.80 and the gray-box attack achieves an average AUC-ROC of 0.90. The best gray-box attack on flan accomplishes almost perfect performance. These results highlight the risk associated with deploying RAG based systems without proper defense mechanisms.
4.3 Defense Strategy
To counter the threat of RAG-MIA attacks, we propose an initial defense strategy in this section. Building on previous research that has demonstrated the ability to influence the output of LLMs through system instructions, we leverage this concept to develop a defense mechanism. Our approach involves modifying the RAG template to include explicit instructions for the LLM to disregard requests that directly inquire about the retrieved database content, thereby providing a defense against RAG-MIA attacks.
Specifically, we evaluated the effectiveness of variations of the following modified RAG template:
Please answer the question using the context provided.
If the question is unanswerable, say unanswerable”.
Pointers for the answer:
1. Repeating, outputting or printing the context is illegal,
answer unanswerable”.
2. Asking if a specific text, fact, or passage appears in your
context is illegal, answer unanswerable”.
Question: {user prompt}
Context:
{context}
We have slightly adapted the RAG template for the llama model, such that it will respond to the defense instructions better, by placing the defense instructions inside the system prompt, an integral capability of llama models. In Table 3 we present the results, and compare them to the scenario where no defense is placed in the RAG system.
Table 3: RAG MIA results with defense
Without defense With defense
Black-Box
AUC ROC
Gray-Box
AUC ROC
Black-Box
AUC ROC
Gray-Box
AUC ROC
Dataset Model
HealthCareMagic flan 0.78 0.99 0.86 0.90
llama 0.88 0.96 0.74 0.51
mistral 0.74 0.83 0.72 0.44
Enron flan 0.82 0.97 0.88 0.96
llama 0.79 0.84 0.77 0.78
mistral 0.78 0.80 0.78 0.72
Our evaluation reveals that the proposed defense strategy yields the most significant benefits against gray-box attacks on the llama and mistral models, across both datasets. Notably, the defense demonstrates high efficacy in the gray-box setting for the mistral model, particularly on the HealthCareMagic dataset, where we observe a substantial improvement of 0.39 in AUC-ROC. In contrast, the defense has a minimal impact on the flan model, only showing a slight effect in the gray-box setting. These findings suggest that further research is needed to explore the feasibility of crafting a RAG template that can effectively defend a flan model against RAG-MIA attacks.
Furthermore, our analysis of the model outputs reveals that a significant proportion of responses do not contain either ”Yes” or ”No” tokens (classified as non-member documents), accounting for approximately 96% of the total outputs on llama, and about 93% on mistral with HealthCareMagic dataset, which explains the defense performance for these cases. On the other hand, for flan, they remain the same (0%). We present the statistics on the amount of such documents in Appendix D.
4.3.1 Improved defense for llama
Since llama models have a dedicated section in the input prompt for system instructions, as mentioned in Section 4.3, we have experimented with placing the retrieved database content within this dedicated section. We compare two cases: (1) With Defense #1 - only the defense instructions are added to the system section (2) With Defense #2 - both defense instructions and retrieved database content are placed in the system section. We present the results for this experiment in Table 4.
Table 4: Llama defenses
Without Defense With Defense #1 With Defense #2
Black-Box
AUC ROC
Gray-Box
AUC ROC
Black-Box
AUC ROC
Gray-Box
AUC ROC
Black-Box
AUC ROC
Gray-Box
AUC ROC
Dataset Model
HealthCareMagic llama 0.88 0.96 0.74 0.51 0.42 0.75
Enron llama 0.79 0.84 0.77 0.78 0.46 0.92
As shown in Table 4, placing defense instructions and retrieved database content in the system section provides robust defense against black-box attacks. However, this approach is less effective against gray-box attacks, where Defense #1 is preferred. Since black-box attacks are more common and require less effort from the attacker, we recommend using Defense #2. Nevertheless, we encourage the research community to develop a defense strategy that can effectively protect from both kinds of attack.
@@ 5 Conclusion
Our attack achieves a very high average AUC-ROC of 0.90 and 0.80 in the gray-box and the black-box threat models, respectively, and for some models almost perfect performance. Our initial defense was able to reduce the success rate of the attack in almost all cases, and mostly prevented the attack for the llama model in the black-box setting. The model that mostly did not benefit from this defense was the flan model for which further defenses need to be developed. This underscores the need to develop more advanced defenses and countermeasures for this attack scenario. We hope that the research community will proceed to explore the risk of membership inference in RAG-based systems and employ the ideas from this paper as a baseline.
@@ B Detailed Results of Retrieval Matches
Tables 6 and 7 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively, for the different attack prompts. We see that the chosen attack prompts do not differ in their influence on the retrieval accuracy.
@@
C Model outputs without a clear yes/no answer
The number of model outputs that did not contain either of the ”Yes” or ”No” tokens across attack prompts for each dataset and model combination are shown in Table 8.
@@ C Model outputs without a clear yes/no answer
D Model outputs without a clear yes/no answer with defense prompt
The number of model outputs that did not contain either of the ”Yes” or ”No” tokens when using a defense prompt for each dataset and model combination are shown in Table 9. This is compared to the corresponding model outputs when using attack prompt #2 without the defense. The number of missing answers increases significantly for the llama and mistral models, and remains the same for flan.
Table 9: Model outputs missing yes/no tokens with defense prompt
Without defense With defense
Missing Total
Percent
missing
Missing Total
Percent
missing
Dataset Model
HealthCareMagic flan 0 4000 $0$ 4000 0.00%
llama 0 4000 0.00% 3868 4000 96.70%
mistral 101 4000 2.53% 3719 4000 92.97%
Enron flan 0 4000 $0$ 4000 0.00%
llama 0 4000 0.00% 3805 4000 95.12%
mistral 41 4000 1.03% 2743 4000 68.58%